Can someone help me on this?


I need to solve 6x + 4(3x - 2) = 6x + 2
thank you, Lisa

Answer by funmath(2933)

Hi Lisa,
I think I've already answered this one twice in the last day or two. If you have already received this answer, I'm not sure what's going on.




I need to solve 6x + 4(3x - 2) = 6x + 2
6x+4(3x-2)=6x+2
6x+4(3x)+4(-2)=6x+2
6x+12x-8=6x+2
18x-8=6x+2
18x-6x-8=6x-6x+2
12x-8=2
12x-8+8=2+8
12x=10
12x/12=10/12
x=5/6
:
Happy Calculating!!!